# Universal Definitions of the Roman Factorial: Introduction to Foundational Functions and the Generalization Process

### v2

@@ 4.5 Second generalization step 2: $\theta(n)$
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot\big{(}n!\big{)}^{1}&,\enspace n\in\mathbb{N}_{0}\\ \eta(n)\cdot\big{[}(-n-1)!\big{]}^{-1}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.5.1)
@@
$\theta(n)=\cfrac{\delta(n)}{|\delta(n)|}=\begin{cases}1&,\enspace n\in\mathbb{N}_{0}\\ -1&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.5.2)
@@
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot\big{(}n!\big{)}^{\theta(n)}&,\enspace n\in\mathbb{N}_{0}\\ \eta(n)\cdot\big{[}(-n-1)!\big{]}^{\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.5.3)
We choose to omit the parentheses enclosing the factorial. It is more concise and clearer to the eye:
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot n!^{\theta(n)}&,\enspace n\in\mathbb{N}_{0}\\ \eta(n)\cdot(-n-1)!^{\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.5.4)
So far, we have introduced two foundational functions into Eq. 4.1.1 in an attempt to rewrite it compactly. However, before introducing another F.F., there is one more adjustment we can make involving the absolute value of $n$.
4.6 Second generalization step 3: $|n|$
@@
$|n|=\begin{cases}n&,\enspace n\geq 0\\ -n&,\enspace n<0\thinspace.\end{cases}$ (Eq. 4.6.1)
We can substitute $|n|$ into Eq. 4.5.4 without any considerations, since the absolute value is a function defined for all reals and thus all integers.
@@
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot|n|!^{\theta(n)}&,\enspace n\in\mathbb{N}_{0}\\ \eta(n)\cdot\big{(}|n|-1\big{)}!^{\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.6.2)
@@
4.7 Second generalization step 4: $\xi^{\prime}(n)$
Let’s investigate the $-1$ that appears only when $n<0$. We can begin by expanding the first case as shown below:
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot\big{(}|n|-0\big{)}!^{\theta(n)}&,\enspace n\in\mathbb{N}_{0}\\ \eta(n)\cdot\big{(}|n|-1\big{)}!^{\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.7.1)
It is evident that we need a function to subtract from $|n|$, which equals 0 for $n\geq 0$ and 1 for $n<0$. Alternatively, we desire a function with the output pattern $[1,0,0]$. This pattern adheres to the F.F. $\xi^{\prime}(n)$:
$\xi^{\prime}(n)=\begin{cases}0&,\enspace n\in\mathbb{N}_{0}\\ 1&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.7.2)
@@ 4.7 Second generalization step 4: $\xi^{\prime}(n)$
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot\big{[}|n|-\xi^{\prime}(n)\big{]}!^{\theta(n)}&,\enspace n\in\mathbb{N}_{0}\\ \eta(n)\cdot\big{[}|n|-\xi^{\prime}(n)\big{]}!^{\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.7.3)
@@
$\lfloor n\rceil!=\eta(n)\cdot\big{[}|n|-\xi^{\prime}(n)\big{]}!^{\theta(n)},\quad n\in\mathbb{Z}\thinspace,$ (Eq. 4.7.4)
@@
$n!=n(n-1)!\thinspace,\quad 0!=1\thinspace,\quad n\in\mathbb{N}\thinspace.$ (Eq. 4.7.5)
@@ 5.2 The function $\Theta(n)$
The first advanced function introduced in this section, denoted by $\Theta(n)$, is desired to output 1 only when $n=0$ and 0 in all other cases. This can be expressed in a piece-wise definition like so:
$\Theta(n)=\begin{cases}1\thinspace,\quad n=0\\ 0\thinspace,\quad n\neq 0\end{cases},\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.2.1)
@@
Therefore, we define $\Theta(n)$ as:
$\Theta(n)=\xi(n)\cdot\xi(-n)\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 5.2.2)
@@
In contrast to previous foundational functions, $\Theta(n)$ appears as a straight line on its plot, except for one point that stands out. In terms of Boolean algebra, $\Theta(n)$ represents the condition of whether $n$ equals zero [^11] More about Boolean algebra in Addendum 9.5..
@@
$\Theta(n)=\begin{cases}\cfrac{\theta(n)+\theta(-n)}{2}\\[11.18028pt] \cfrac{\xi(n)+\xi(-n)}{2}\\[11.18028pt] \cfrac{1+\theta(n)\cdot\theta(-n)}{2}\\[11.18028pt] \cfrac{1+\theta(-|n|)}{2}\\[11.18028pt] \xi(-|n|)\end{cases},\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.2.3)
@@
As an exercise, one can delve deeper into understanding these relationships by proving them. This can be accomplished through mathematical proofs or by visualizing the individual functions’ graphs [^12] All graphs in this paper were made with Wolfram Mathematica. and their combinations.
@@ 5.3 The function $Q(n)$
$Q(n)=\begin{cases}\theta(n)\thinspace,&n>0\\ 0\thinspace,&n=0\\ \theta(n)\thinspace,&n<0\end{cases}\quad,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.3.2)
@@
$Q(n)=\begin{cases}\theta(n)\thinspace,&n\neq 0\\ 0\thinspace,&n=0\end{cases}\quad,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.3.3)
@@
$Q(n)=\begin{cases}\theta(n)-0\thinspace,&n\neq 0\\ \theta(n)-1\thinspace,&n=0\end{cases}\quad,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.3.4)
@@
$Q(n)=\theta(n)-\Theta(n)\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 5.3.5)
@@
$Q(n)=\begin{cases}\cfrac{\theta(n)-\theta(-n)}{2}\\[8.68027pt] \xi(n)-\xi(-n)\\[6.18027pt] \xi^{\prime}(-n)-\xi^{\prime}(n)\end{cases},\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.3.6)
@@ 5.4 The function $Q^{\prime}(n)$
The function $Q^{\prime}(n)$ is akin to $\Theta(n)$, but with reversed criteria. In essence, $Q^{\prime}(n)$ yields 0 only when $n=0$ and 1 for all other $n$. Thus, the piece-wise definition of $Q^{\prime}(n)$ is as follows:
$Q^{\prime}(n)=\begin{cases}0\thinspace,&n=0\\ 1\thinspace,&n\neq 0\end{cases}\quad,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.4.1)
An easy approach to achieve this goal using the F.F. is to take the absolute value of $Q(n)$. Since $Q(n)$ has the output pattern $[-1,0,1]$, we can notice that $|Q(n)|$ will have the pattern $[1,0,1]$. That is our desired set of outputs, so we can define $Q^{\prime}(n)$ as:
$Q^{\prime}(n)=|Q(n)|\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 5.4.2)
@@
$Q^{\prime}(n)=\begin{cases}1-\Theta(n)\\[8.68027pt] \cfrac{1-\theta(n)\cdot\theta(-n)}{2}\\[8.68027pt] \xi^{\prime}(-|n|)\end{cases},\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.4.3)
@@ 5.5 The function $\Psi(n)$
The next function we’ll introduce in this paper is $\Psi(n)$. We need this function to be essentially $n$ for all $n\neq 0$, and 1 for $n=0$. Its requirements can be expressed in the following piece-wise definition:
$\Psi(n)=\begin{cases}n\thinspace,&n\neq 0\\ 1\thinspace,&n=0\end{cases}\quad,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.5.1)
@@
$\Psi(n)=n+\Theta(n)\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 5.5.2)
@@ 5.6 The function $\Phi(n)$
$\Phi(n)=\begin{cases}1\thinspace,&n\geq 0\\ n\thinspace,&n<0\end{cases}\quad,\enspace n\in\mathbb{R}\thinspace.$ (Eq. 5.6.1)
@@
$\Phi(n)=n^{\xi^{\prime}(n)}\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 5.6.2)
This works well, except for the case when $n=0$: $\Phi(0)=0^{0}$, which is problematic because $0^{0}$ is undefined [^13] The indeterminate form $0^{0}$ does not have a universally agreed-upon value, although it is given the value 1 in a lot of contexts [3]. More about $0^{0}$ in Addendum 9.6.. Fortunately we can solve this problem by using $\Psi(n)$. We end up with this universal definition of $\Phi(n)$:
$\Phi(n)={\Psi(n)}^{\xi^{\prime}(n)}\thinspace,\quad n\in\mathbb{R}\thinspace.$ (Eq. 5.6.3)
@@ 5.7 Synopsis
Conclusively, this section presents another set of 5 foundational functions. They were all expressed in terms of other F.F., and along with those in Section 3 they constitute a set of Boolean-like functions that are instrumental to this paper.
@@
For example, the output pattern of $\xi^{\prime}(n)$ is $[1,0,0]$. That can be read as the number $100_{2}$, which is equal to 4 in base 10. Many F.F. exhibit this behavior, but those with outputs like $-1$ or $n$ are excluded.
In this way, we can construct a binary table. Tbl. 5.7.2 has all 3-digit binary numbers:
@@
The completion of this table helped define some of these functions. After discovering a few of them, there was motivation to finish Tbl. 5.7.2. However, it’s important to note that the binary representation doesn’t have any special meaning; it just helps organize the table.
Since there are many new functions introduced in this paper, it’s easy to get them mixed up. To be reminded of what each function does, refer to Eq. 5.7.1. That list will ensure that everything in the paper is clear and easy to understand.
@@
6 Non-recursive Roman factorial

### v7

@@ 4.5 Roman-like definition step 2: $\theta(n)$
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot\big{(}n!\big{)}^{1}&,\enspace n\in\mathbb{Z}^{+}_{0}\\[3.68027pt] \eta(n)\cdot\big{[}(-n-1)!\big{]}^{-1}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.5.1)
@@
$\theta(n)=\cfrac{\delta(n)}{|\delta(n)|}=\begin{cases}1&,\enspace n\in\mathbb{Z}^{+}_{0}\\ -1&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 3.3.2)
@@
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot\big{(}n!\big{)}^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{+}_{0}\\[3.68027pt] \eta(n)\cdot\big{[}(-n-1)!\big{]}^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.5.2)
We will omit the parentheses enclosing the factorial. Now, it is more concise and clear to the eye:
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot n!^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{+}_{0}\\[3.68027pt] \eta(n)\cdot(-n-1)!^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.5.3)
So far, we have introduced two foundational functions into Eq. 2.2.1 in an attempt to rewrite it compactly. Before introducing another F.F. however, there is one more adjustment we can make involving the absolute value of $n$.
4.6 Roman-like definition step 3: $|n|$
@@
$|n|=\begin{cases}n&,\enspace n\geq 0\\ -n&,\enspace n<0\,.\end{cases}$ (Eq. 4.6.1)
We can substitute $|n|$ into Eq. 4.5.3 without any further consideration, since the absolute value is a function defined for all reals and thus all integers.
@@
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot|n|!^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{+}_{0}\\[3.68027pt] \eta(n)\cdot\big{(}|n|-1\big{)}!^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.6.2)
@@
4.7 Roman-like definition step 4: $\xi^{\prime}(n)$
Let’s investigate the $-1$ that appears only when $n\in\mathbb{Z}^{-}$. We can begin by expanding the first case as shown below:
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot\big{(}|n|-0\big{)}!^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{+}_{0}\\[3.68027pt] \eta(n)\cdot\big{(}|n|-1\big{)}!^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.7.1)
It is evident that we need a function to subtract from $|n|$, which equals 0 for $n\in\mathbb{Z}^{+}_{0}$ and 1 for $n\in\mathbb{Z}^{-}$. Alternatively, we desire a function with the output pattern $[1,0,0]$. This pattern adheres to the F.F. $\xi^{\prime}(n)$:
$\xi^{\prime}(n)=\frac{1-\theta(n)}{2}=\begin{cases}0&,\enspace n\in\mathbb{Z}^{+}_{0}\\ 1&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 3.5.1)
@@ 4.7 Roman-like definition step 4: $\xi^{\prime}(n)$
$\lfloor n\rceil!=\begin{cases}\eta(n)\cdot\big{(}|n|-\xi^{\prime}(n)\big{)}!^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{+}_{0}\\[3.68027pt] \eta(n)\cdot\big{(}|n|-\xi^{\prime}(n)\big{)}!^{\,\theta(n)}&,\enspace n\in\mathbb{Z}^{-}.\end{cases}$ (Eq. 4.7.2)
@@
$\lfloor n\rceil!=\eta(n)\cdot\big{(}|n|-\xi^{\prime}(n)\big{)}!^{\,\theta(n)}\,,\quad n\in\mathbb{Z}\,,$ (Eq. 4.7.3)
@@
$n!=n(n-1)!\,,\quad 0!=1\,,\quad n\in\mathbb{Z}^{+}.$ (Eq. 2.1.3)
@@ 5.2 The function $\Theta(n)$
The first advanced function [^14] These advanced functions are not more difficult to understand, but they are more intricate, as they consist of many previously defined foundational functions. introduced in this section, denoted by $\Theta(n)$, is desired to output 1 only when $n=0$ and 0 in all other cases. This can be expressed in a piece-wise definition like so:
$\Theta(n)=\begin{cases}1\,,\quad n=0\\ 0\,,\quad n\neq 0\end{cases},\enspace n\in\mathbb{R}\,.$ (Eq. 5.2.1)
@@
Therefore, we define $\Theta(n)$ universally as:
$\Theta(n)=\xi(n)\cdot\xi(-n)\,,\quad n\in\mathbb{R}\,.$ (Eq. 5.2.2)
@@
In contrast to previous foundational functions, $\Theta(n)$ appears as a straight line on its plot, except for one point that stands out. In terms of Boolean algebra, $\Theta(n)$ represents the condition of whether $n$ equals zero.
@@
$\Theta(n)=\begin{cases}\cfrac{\theta(n)+\theta(-n)}{2}\\[11.18028pt] \cfrac{\xi(n)+\xi(-n)}{2}\\[11.18028pt] \cfrac{1+\theta(n)\cdot\theta(-n)}{2}\\[11.18028pt] \cfrac{1+\theta\big{(}-|n|\big{)}}{2}\\[8.68027pt] \xi\big{(}-|n|\big{)}\end{cases},\enspace n\in\mathbb{R}\,.$ (Eq. 5.2.3)
@@
As an exercise, one can delve deeper into understanding these relationships by proving them. This can be accomplished through mathematical proofs or by visualizing the individual functions’ graphs [^15] All graphs in this paper were made with Wolfram Mathematica. For help with the relevant code, contact information is at Section 8. and their combinations.
@@ 5.3 The function $Q(n)$
$Q(n)=\begin{cases}\theta(n)\,,&n>0\\ 0\,,&n=0\\ \theta(n)\,,&n<0\end{cases}\quad,\enspace n\in\mathbb{R}\,.$ (Eq. 5.3.2)
@@
$Q(n)=\begin{cases}\theta(n)\,,&n\neq 0\\[1.18027pt] 0\,,&n=0\end{cases}\quad,\enspace n\in\mathbb{R}\,.$ (Eq. 5.3.3)
@@
$Q(n)=\begin{cases}\theta(n)-0\,,&n\neq 0\\[1.18027pt] \theta(n)-1\,,&n=0\end{cases}\quad,\enspace n\in\mathbb{R}\,.$ (Eq. 5.3.4)
@@
$Q(n)=\theta(n)-\Theta(n)\,,\quad n\in\mathbb{R}\,.$ (Eq. 5.3.5)
@@
$Q(n)=\begin{cases}\cfrac{\theta(n)-\theta(-n)}{2}\\[8.68027pt] \xi(n)-\xi(-n)\\[6.18027pt] \xi^{\prime}(-n)-\xi^{\prime}(n)\end{cases},\enspace n\in\mathbb{R}\,.$ (Eq. 5.3.6)
@@ 5.4 The function $Q^{\prime}(n)$
The function $Q^{\prime}(n)$ is akin to $\Theta(n)$, but with reversed criteria. In essence, $Q^{\prime}(n)$ yields 0 only when $n=0$ and 1 for $n\neq 0$. Thus, the piece-wise definition of $Q^{\prime}(n)$ is as follows:
$Q^{\prime}(n)=\begin{cases}0\,,&n=0\\ 1\,,&n\neq 0\end{cases}\quad,\enspace n\in\mathbb{R}\,.$ (Eq. 5.4.1)
An easy approach to achieve this goal using the F.F. is to take the absolute value of $Q(n)$. Since $Q(n)$ has the output pattern $[-1,0,1]$, we see that $|Q(n)|$ will have the pattern $[1,0,1]$. That is our desired set of outputs, so we can define $Q^{\prime}(n)$ as the absolute value of $Q(n)$:
$Q^{\prime}(n)=|Q(n)|\,,\quad n\in\mathbb{R}\,.$ (Eq. 5.4.2)
@@
$Q^{\prime}(n)=\begin{cases}1-\Theta(n)\\[8.68027pt] \cfrac{1-\theta(n)\cdot\theta(-n)}{2}\\[8.68027pt] \xi^{\prime}\big{(}-|n|\big{)}\end{cases},\enspace n\in\mathbb{R}\,.$ (Eq. 5.4.3)
@@ 5.5 The function $\Psi(n)$
The next function we’ll introduce in this paper is $\Psi(n)$. We need this function to be equal to $n$ for all $n\neq 0$, and 1 for $n=0$. Its requirements can be expressed in the following piece-wise definition:
$\Psi(n)=\begin{cases}n\,,&n\neq 0\\ 1\,,&n=0\end{cases}\quad,\enspace n\in\mathbb{R}\,.$ (Eq. 5.5.1)
@@
$\Psi(n)=n+\Theta(n)\,,\quad n\in\mathbb{R}\,.$ (Eq. 5.5.2)
@@ 5.6 The function $\Phi(n)$
$\Phi(n)=\begin{cases}1\,,&n\geq 0\\ n\,,&n<0\end{cases}\quad,\enspace n\in\mathbb{R}\,.$ (Eq. 5.6.1)
@@
$\Phi(n)=n^{\xi^{\prime}(n)}\,,\quad n\in\mathbb{R}\,.$ (Eq. 5.6.2)
This works well, except for the case when $n=0$: $\Phi(0)=0^{0}$, which is problematic because $0^{0}$ is undefined [^16] The indeterminate form $0^{0}$ does not have a universally agreed-upon value, although it is given the value 1 in a lot of contexts [0tothe0]. More about $0^{0}$ in Addendum 9.6.. Fortunately we can solve this problem by using $\Psi(n)$. We end up with this universal definition of $\Phi(n)$:
$\Phi(n)={\Psi(n)}^{\xi^{\prime}(n)}\,,\quad n\in\mathbb{R}\,.$ (Eq. 5.6.3)
@@ 5.7 Synopsis
In short, this section is about presenting another set of 5 foundational functions. They were all expressed in terms of other F.F., and along with those in Section 3 they constitute a set of Boolean-like functions that are instrumental to this paper.
@@
$\delta(n)$ $=\lfloor n\rfloor+0.5$ [,+,+[-,+,+][ - , + , + ] (Eq. 3.2.3)
$\theta(n)$ $=\cfrac{\delta(n)}{\left|\delta(n)\right|}$ $[-1,1,1]$ (Eq. 3.3.2)
$\xi(n)$ $=\cfrac{1+\theta(n)}{2}$ $[0,1,1]$ (Eq. 3.4.2)
$\xi^{\prime}(n)$ $=\cfrac{1-\theta(n)}{2}$ $[1,0,0]$ (Eq. 3.5.1)
@@
$\Theta(n)$ $=\xi(n)\cdot\xi(-n)$ $[0,1,0]$ (Eq. 5.2.2)
$Q(n)$ $=\theta(n)-\Theta(n)$ $[-1,0,1]$ (Eq. 5.3.5)
$Q^{\prime}(n)$ $=1-\Theta(n)$ $[1,0,1]$ (Eq. 5.4.2)
$\Psi(n)$ $=n+\Theta(n)$ $[n,1,n]$ (Eq. 5.5.2)
$\Phi(n)$ $={\Psi(n)}^{\xi^{\prime}(n)}$ $[n,1,1]$ (Eq. 5.6.3)
@@
For example, the output pattern of $\xi^{\prime}(n)$ is $[1,0,0]$. That can be read as the binary number $100_{2}$, which is equal to 4 in base 10. Many F.F. exhibit this behavior, but those with outputs like $-1$ or $n$ are excluded.
In this way, we can construct a binary table. The Tbl. 5.7.2 has all 3-digit binary numbers:
@@
The completion of this table helped define some of these functions. After discovering a few of them, there was motivation to finish Tbl. 5.7.2. However, it’s important to note that the binary representation doesn’t have any applications for this paper; it just helps organize the table.
Since there are many new functions introduced in this paper, it’s easy to get them mixed up. Refer to the previous equation list to be reminded of what each function does. That list will ensure that everything in the paper is clear and easy to understand.
@@
6 $\prod$-product generalizations of $\lfloor n\rceil!$